OpenStudy (anonymous):

Is the derivative of 2sec(x) = 2(sec(x)tan(x)) or 2sec(x)tan(x)?

OpenStudy (anonymous):

2 sec(x)tan(x)

OpenStudy (anonymous):

i feel strongly that they are the same thing

OpenStudy (anonymous):

oh i didnt see the extra bracket. nvm

OpenStudy (anonymous):

\[\ln (x/2 + \sqrt{x^2-4}/2) = \ln (x+\sqrt{x^2-4})?\]

OpenStudy (anonymous):

I am not sure how this is true

OpenStudy (anonymous):

How does the 2 come off the bottom?

OpenStudy (anonymous):

is the other 2 under the square root sign or not?

OpenStudy (anonymous):

No the x and the sqrt(x^2-4) are both being divided by two, but for some reason the book loses the 2 on the bottom.

OpenStudy (anonymous):

yeah they just didvide through by 2. If you nice the 2 under the single x is gone too

OpenStudy (anonymous):

Explain this, I am just a bit confused.

OpenStudy (anonymous):

The equation above was just the final answer, not sure how it was achieved.